OpenStudy (lilsis76):

Geometric Sequences and Series: how can i express each repeating decimal as a fraction? .123.... i know the 1 is in the tenths, 23 is in the 1000 or 10,000.

OpenStudy (anonymous):

\[=\frac{ 1 }{ 10 }+\frac{ 2 }{ 10^{2} }+\frac{ 3 }{ 10^{3} }...\]

OpenStudy (lilsis76):

okay, is that all i would have to do?

OpenStudy (anonymous):

That answers the question you posted :)